# Some configuration results for area-minimizing cones

### v4

@@ Abstract.
We discover some configuration results for area-minimizing cones. In particular, given any closed minimal submanifold in some Euclidean sphere, every cone over the minimal product of sufficiently many copies of the submanifold turns out to be area-minimizing.
@@
Area-minimizing cone, Lawlor criterion
1991 Mathematics Subject Classification:
Primary 28A75, Secondary 53C38
@@ 1. Introduction
When $L$ is an embedded closed smooth submanifold, the cone $C(L)$ is singular only at the origin and we say it is a regular cone. In this paper, we search for regular area-minimizing cones and end up with very general configuration results.
@@
Apparently, the minimal product can apply to multiple inputs inductively. Along this line, Lawlor (Theorem 5.1.1, [Law]) completely settled down $(\star)$ when each input $L_{i}$ is $\mathbb{S}^{N_{i}}$, with $i=1,2,\cdots,n$ and $N_{i}\in\mathbb{Z}_{+}$. In particular, whenever $\Sigma_{i=1}^{n}N_{i}>7$, the corresponding cone is area-minimizing. In recent work [TZ], cones of dimension no less than 37 over minimal products of the category of typical minimal submanifolds associated to isoparametric foliations of spheres are confirmed area-minimizing. From these, one can observe that the dimension plays an important role for $(\star)$ and that some minimal submanifolds in spheres themselves cannot span area-minimizing cones but their products do.
@@
The result tells in particular that every closed minimal submanifold with no other assumptions at all can be used as a brick to construct area-minimizing cones in different combinatorial ways. The only payment is that the resulting area-minimizing cone may have high dimension.
Moreover, we establish the following using the terminology from the set theory.
@@
Usually, it is not easy to construct regular area-minimizing cones before. But Theorem 1.2 shows that there are a plethora of such kind cones and the cardinality is decided by closed minimal submanifolds in Euclidean spheres.
The main tool is Lawlor’s criterion for a regular cone to be area-minimizing. To be self-contained we briefly explain his method in §2. Some useful information about minimal product will be given in §3. Finally, the proof of Theorems 1.1 and 1.2 and more subtle results (also interesting) by applying Lawlor’s criterion are collected in §4.
Acknowledgment. The author wishes to thank Professor Blaine Lawson for drawing his attention to the powerful work [Law] by Professor Gary Lawlor.
@@ 2. Lawlor’s curvature criterion
Area-minimizing hypercone is a key ingredient to solve the famous Bernstein problem [B] about entire minimal graph of codimension one, see [fle, de2, A]. It was Jim Simons who found first non-trivial stable minimal hypercones $C(S^{k}\times S^{k})$ with $k\geq 3$ in [JS] and suspected that they are area-minimizing. In the following year, breakthrough was made in [BdGG] which confirmed those Simons cones’ area-minimality and based on this counterexamples to the Bernstein problem were found over graph domain $\mathbb{R}^{N}$ when $N\geq 8$.
In [HS] a beautiful characterization of area-minimizing hypercones was established. In each side of an area-minimizing hypercone, there is a unique dilation-invariant foliation by minimal hypersurfaces. Therefore, the unit norm vector fields lead to a foliation by their integral curves on $\mathbb{R}^{N}\sim 0$, which is also invariant under dilation. It naturally induces a dual co-degree one calibration form (see [HL]) and anl area-nonincreasing projection (for surfaces of codimension one) to the hypercone.
For a given regular cone, how to determine whether it is area-minimizing or not? The most effective way is to employ the theory of calibration. However, in general, it is not easy to directly find a desired calibration with mild singularity. Lawlor successfully developed a sufficient criterion (sometimes also necessary) looking for area-nonincreasing projection structure in suitable angular neighborhood of the minimal cone instead of in the whole space $\mathbb{R}^{N}$. If the boundary of the neighborhood is mapped to the origin under the projection, then one can send everything outside the neighborhood to the origin. In this way an area-nonincreasing projection can be globally defined. However, its dual calibration has quite “large” singularity.
To understand Lawlor’s criterion, let us recall the following for $k<N$.
@@
Suppose $L^{k}\subset\mathbb{S}^{N}$ is a closed submanifold and $x\in L$. Let $S_{x}$ denote the set of all unit normals of $L$ at $x$. Then the normal radius of $L$ (or $C(L)$) at $x$ is
$R_{x}\triangleq\sup\left\{\,s\in\mathbb{R}_{+}\,:\,\exp_{x}(sv_{x})\not\in L\,\text{ for all }v_{x}\in S_{x}\right\}$
where $\exp_{x}(sv_{x})=(\cos s)x+(\sin s)v_{x}$ is a normal geodsic. The quantity $R(L)\triangleq\inf\{\,R_{x}\,:\,x\in L\,\}$ is called normal radius of $L$ (or $C(L)$).
To establish area-minimizing projection, Lawlor considered foliation generated by rotations of a suitable curve $\gamma_{x}$ and dilations in each normal wedge
$W_{x}\left(\frac{R_{x}}{2}\right)=C\left(\exp_{x}\left(\left[0,\frac{R_{x}}{2}\right]\cdot S_{x}\right)\right).$
@@
with $r(0)=\|x\|=1$. Here $\textbf{h}_{ij}^{v}$ means the second fundamental form at $x$ with respect to $v$. To get narrowest curve $\gamma_{x}$, one needs to require that equality holds everywhere in (2.1) and wants to have a smallest $\theta_{0}$ such that $\lim_{\theta\uparrow\theta_{0}}r(\theta)=\infty$. As a result, the boundary of $W_{x}\left(\theta_{0}\right)$ is projected to the origin and global area-decreasing projection gets set up. Such $\theta_{0}$, if existed, is called vanishing angle of $C(L)$. Note that there is no focal point of $L$ in the neighborhood $\bigcup_{x\in L}W_{x}\left(\theta_{0}\right)$ since otherwise that focal point leads to a strictly negative expression under the square root symbol in (2.1).
@@
By the control
@@
where $t=\tan\theta$, $\alpha=\sup_{x\in L,v\in S_{x}}\|\textbf{h}_{ij}^{v}\|$ and
@@
Lawlor derived two kinds of vanishing angles by employing (2.2) to replace the infimum term in (2.1)
@@
So, if either $\theta_{F}$ or $\theta_{c}$ exists of finite value, then consequently vanishing angle $\theta_{0}$ must exist. For details readers are referred to [Law]. In this paper we will focus on the following nice property.
@@
In fact a ratio type estimate on $\theta_{F}$ can also be established.
@@
In this paper we will in particular consider the following subclass.
@@
We will show that the situation in Theorem 1.1 brings us Type-c cones (also Type-F cones). Besides, some considerations regarding $(\star)$ will be made in §4 (and results work for Type-F as well).
@@ 3. Minimal product
This structure should be a very classical result. There have been many specific situations where minimal product has been employed. For all $L_{i}$ being full spheres, the structure produces classical Clifford type minimal surfaces as already mentioned in §1. When these $L_{i}$ are taken to be R-spaces, [OS] studied whether cones over minimal products of them are area-minimizing. For general minimal submanifolds as inputs, see [X] for a proof through the Takahashi Theorem. In 2016 (arXiv chronology), it was rediscovered independently by [TZ] and [CH] using different methods. Reader may choose either familiar way to understand the structure.
3.1. Estimate on $\alpha^{2}$ for minimal product.
For simplicity, let us fix $n=2$ in the paper and all related results for $n\geq 3$ can be gained by induction. In [TZ], information about the second fundamental form of minimal product has been clearly figured out. Let $k_{1},k_{2}$ be the dimension of $L_{1},L_{2}$ respectively. Take $x_{1}\in L_{1}$ and $x_{2}\in L_{2}$. Together with $\eta_{0}=(\lambda_{2}x_{1},-\lambda_{1}x_{2})$, orthonormal bases $\{\sigma_{1},\cdots,\sigma_{N_{1}-k_{1}}\}$ and $\{\tau_{1},\cdots,\tau_{N_{2}-k_{2}}\}$ of $T_{x_{1}}^{\perp}L_{1}$ and $T_{x_{2}}^{\perp}L_{2}$ lead to an orthonormal basis
@@
3.2. Normal radius of minimal product
Let $R_{1},\,R_{2},\,R$ be the normal radii of $L_{1},\ L_{2}$ and $L_{1}\times L_{2}$ in corresponding Euclidean spheres respectively.
For $(\lambda_{1}x_{1},\lambda_{2}x_{2})$ in $L_{1}\times L_{2}$, the combination
@@ 3.2. Normal radius of minimal product
where $\xi_{1}$, $\xi_{2}$ are unit normals of $L_{1}$ and $L_{2}$ at $x_{1}$ and $x_{2}$ respectively, and $a_{1}^{2}+a_{2}^{2}+a_{0}^{2}=1$, can exhaust all unit normals of $L_{1}\times L_{2}$ at $p$. Then, in $\mathbb{S}^{N_{1}+N_{2}+1}$, each unit normal induces a normal geodesic
@@
So, to search for the value of $R$ one should focus on Case II. Note that it is always true that
@@
Hence $\sin R\neq 0$. In Case II, following the discussion of [JCX] in fact one can show that no matter what value of $\phi_{1}$ is in (3.4), the triple $(a_{0},a_{1},a_{2})$ can be solved to make (3.4) hold. Namely,
(3.8) $-a_{0}\lambda_{2}=\lambda_{1}\frac{\cos R-\cos\phi_{1}}{\sin R},$
(3.9) $a_{0}\lambda_{1}=\lambda_{2}\frac{\cos R-1}{\sin R},\ \ \$
$a_{2}=0.\ \ \ \ \ \ \ \ \ \ \ \$
These can be read off directly from (3.4) and (3.5). However, the reverse logic will be more important to us, since what we want to solve is this:
Given $\phi_{1}=R_{1}$ for some $x_{1}$ and direction $\xi_{1}$ realizing normal radius $R_{1}$ of $L_{1}$ in $\mathbb{S}^{N_{1}}$, can we get $a_{0},\,a_{1}$ (with $a_{0}^{2}+a_{1}^{2}=1$) and some number $R$ to make (3.4) and (3.5) hold?
If affirmative, we know that the equality of (3.7) stands. This becomes more clear from its solution
(3.10) $\cos R=\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2},$
@@
We remark that the above argument works perfectly for genuine submanifolds of spheres. However, when $L_{1}$ (or $L_{2}$ or both) is a full sphere, the normal radius $R_{1}$ (or $R_{2}$ or both) cannot be well defined as in Definition 2.1. For our purpose, we always define the normal radius of a full sphere $\mathbb{S}^{N_{1}}$ in itself to be $\pi$. The reason is the following. It can be realized as a totally geodesic sphere in $\mathbb{S}^{N_{1}+\ell}$ for $\ell>0$ and now $\mathbb{S}^{N_{1}}$ has normal radius $\pi$. So the above argument works for the minimal product and the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ can be completely dropped as they contribute nothing (let $\xi$ be a unit normal vector to the minimal product at some point and if $\xi$ has a nonzero part coming from the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ then the geodesic with initial velocity vector $\xi$ in $\mathbb{S}^{N_{1}+\ell+N_{2}}$ will run at least length $\pi$ to return to the minimal product and hence this will not realize the normal radius as the normal radius is strictly less than $\pi$ for the minimal product - not a totally geodesic sphere, actually no more than $\frac{\pi}{2}$ by Lemma 4.7). For example, the normal radius of minimal product $\mathbb{S}^{N_{1}}\times\mathbb{S}^{N_{2}}$ in $\mathbb{S}^{N_{1}+N_{2}+1}$ is realized by $\eta_{0}$ and can reach $\frac{\pi}{2}$ only when $N_{1}=N_{2}$ according to (3.10), also see the beautiful uniform formula in [Law] for normal radius of minimal product among full spheres.
As a result, with the above understanding we arrive at the following.
Lemma 3.2.
Proof.
This is exactly given by the equality situation in Case II and take the larger value of $\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}$ and $\lambda_{1}^{2}+\lambda_{2}^{2}\cos R_{2}$. ∎
Remark 3.3.
Note that the result is in fact valid when $\lambda_{1}$ and $\lambda_{2}$ are just two nonzero real numbers with $a_{1}^{2}+a_{2}^{2}=1$. However, we will restrict ourselves in the minimal product construction unless otherwise specified.
@@
Not as in §3.1, the formula here does not need $\{L_{i}\}$ to be minimal but depends on the constant multipliers $\{\lambda_{i}\}$ in the construction of the product.
Remark 3.7.
Statement formulated independently in [JCX].

### v5

@@ Abstract.
We discover some very general configuration results for constructing area-minimizing cones. In particular, given any closed minimal submanifold in some Euclidean sphere, every cone over the minimal product of sufficiently many copies of the submanifold turns out to be area-minimizing; meanwhile every cone over the minimal product of the submanifold and a round sphere of sufficiently large dimension is also area-minimizing. Here no additional geometric assumption (e.g. on isometry group or second fundamental form) is required. Moreover, we establish that the category of regular area-minimizing cones in Euclidean spaces and that of closed minimal submanifolds in Euclidean spheres share the same cardinality.
@@
Regular area-minimizing cone, Lawlor criterion
@@ 1. Introduction
When $L$ is an embedded closed smooth submanifold, the cone $C(L)$ is smooth away from the origin and we say it is a regular cone. In this paper, we search for regular area-minimizing cones and end up with very general configuration results.
@@
Apparently, the minimal product can apply to multiple inputs inductively. Along this line, Lawlor (Theorem 5.1.1, [Law]) completely settled down $(\star)$ when each input $L_{i}$ is $\mathbb{S}^{N_{i}}$, with $i=1,2,\cdots,n$ and $N_{i}\in\mathbb{Z}_{+}$. In particular, whenever $\Sigma_{i=1}^{n}N_{i}>7$, the corresponding cone is area-minimizing. In recent work [TZ], cones of dimension no less than 37 over minimal products of the category of typical minimal submanifolds associated to isoparametric foliations of spheres are confirmed area-minimizing. From these, one can observe that the dimension plays an important role for $(\star)$ and that some minimal submanifolds in spheres themselves cannot span area-minimizing cones but their minimal products do.
@@
Remark 1.2.
Every closed minimal submanifold with no other assumptions at all can be used as a brick to construct area-minimizing cones in different combinatorial ways. The only payment is that the resulting area-minimizing cone may have high dimension.
Another neat result is the following.
@@
Remark 1.4.
In fact there exists integer $d_{0}$ such that the cone over the minimal product $L^{k}\times\mathbb{S}^{d_{1}}\times\cdots\mathbb{S}^{d_{\ell}}$ is area-minimizing whenever $d=\sum_{i=1}^{\ell}d_{i}\geq d_{0}$. If $L$ is a minimal isoparametric hypersurface or a focal submanifold, then this falls into [TZ] and $k+d\geq 36$ would be sufficient for the area-minimality.
Prior to this paper, all known regular area-minimizing cones are tightly related to either group actions or the theory of isoparametric foliations on spheres. The above theorems in particular break these restrictions. Moreover, a corollary of the theorems from the point of set theory is established as follow.
Corollary 1.5.
@@
The main tool is Lawlor’s curvature criterion for a regular cone to be area-minimizing. To be self-contained we briefly explain his method in §2. Some useful information about minimal product will be given in §3. Finally, the proofs of the theorems and corollary and more subtle results partially answering $(\star)$ and complementing the theorems are collected in §4.
Acknowledgments. The author wishes to thank Professor Blaine Lawson for drawing his attention to the powerful work [Law] of Professor Gary Lawlor. This work was sponsored in part by NSFC (Grant Nos. 12022109, 11971352).
@@ 2. Lawlor’s curvature criterion
Area-minimizing hypercone is a key ingredient to solve the famous Bernstein problem [B] about entire minimal graph of codimension one, see [fle, de2, A]. It was Jim Simons who found first non-trivial stable minimal hypercones $C(S^{k}\times S^{k})$ with $k\geq 3$ in [JS] and suspected that they are area-minimizing. In the following year, breakthrough was made in [BdGG] which confirmed those Simons cones’ area-minimality and based on this discovery counterexamples of non-planar minimal graphs to the Bernstein problem were found over domain $\mathbb{R}^{N}$ when $N\geq 8$.
In [HS] a beautiful characterization of area-minimizing hypercones was established. In each side of an area-minimizing hypercone, there is a unique dilation-invariant foliation by minimal hypersurfaces. Therefore, the unit norm vector fields lead to a foliation by their integral curves on $\mathbb{R}^{N+1}\sim 0$, which is also invariant under dilation. It naturally induces a dual co-degree one calibration form (see [HL]) and an area-nonincreasing projection (for surfaces of codimension one) to the hypercone in $\mathbb{R}^{N+1}\sim 0$.
For a given regular minimal cone, how to determine whether it is area-minimizing or not? The most effective way is to employ the theory of calibration. However, in general, it is not easy to directly find a desired calibration with mild singularity. Lawlor successfully developed a sufficient criterion (sometimes also necessary) looking for area-nonincreasing projection structure in suitable angular neighborhood of the minimal cone instead of in the whole space $\mathbb{R}^{N}$. If the boundary of the neighborhood is mapped to the origin under the projection, then one can send everything outside the neighborhood to the origin. In this way an area-nonincreasing projection can be globally defined.
To understand Lawlor’s criterion, let us recall the following with $k<N$.
@@
Let $L^{k}\subset\mathbb{S}^{N}$ be a closed embedded submanifold, $x\in L$ and $S_{x}$ denote the set of all unit normals of $L$ at $x$. Then the normal radius of $L$ at $x$ is
$R_{x}\triangleq\sup\left\{\,t\in\mathbb{R}_{+}\,:\,\exp_{x}(sv_{x})\not\in L\,\text{ for all }v_{x}\in S_{x}\text{ and }s\in(0,t)\right\}$
where $\exp_{x}(sv_{x})=(\cos s)\,x+(\sin s)\,v_{x}$ forms a normal geodesic along $v_{x}$. The quantity $R(L)\triangleq\inf\{\,R_{x}\,:\,x\in L\,\}$ is called normal radius of $L$.
For gaining area-minimizing projection, Lawlor considered in each normal wedge
$W_{x}\left(\frac{R_{x}}{2}\right)=C\left(\exp_{x}\left(\left[0,\frac{R_{x}}{2}\right]\cdot S_{x}\right)\right)$
@@
with $r(0)=\|x\|=1$. Here $\textbf{h}_{ij}^{v}$ means the second fundamental form at $x$ with respect to $v$. To get narrowest curve $\gamma_{x}$, one needs to require that equality holds everywhere in (2.1) and wants to have the smallest $\theta_{0}$ such that $\lim_{\theta\uparrow\theta_{0}}r(\theta)=\infty$. As a result, the boundary of the normal wedge $W_{x}\left(\theta_{0}\right)$ is projected to the origin and global area-decreasing projection gets set up. Such $\theta_{0}$, if existed, is called vanishing angle of $C(L)$. Note that there is no focal point of $L$ in the neighborhood $\bigcup_{x\in L}W_{x}\left(\theta_{0}\right)$ since otherwise that focal point leads to a strictly negative expression under the square root symbol in (2.1).
@@
Where does curvature play a role? How to check whether a given cone satisfies $\theta_{0}\leq\frac{R(L)}{2}$? In practice, Lawlor employed
@@
where $t=\tan\theta$, $\alpha=\sup_{x\in L,v\in S_{x}}\|\textbf{h}_{ij}^{v}\|$ is the curvature term and
@@
to replace the infimum term in (2.1) deriving two kinds of vanishing angles
@@
So, if either $\theta_{F}$ or $\theta_{c}$ exists of finite value, then consequently vanishing angle $\theta_{0}$ must exist. For details readers are referred to [Law].
An important property about $\theta_{c}$ is the following, which provides an effective control on $\theta_{c}$.
@@
In fact a ratio type control on $\theta_{F}$ can also be established.
@@
We will show that the situations in Theorems 1.1 and 1.3 bring us Type-c cones. Besides, some considerations regarding $(\star)$ will be made in §4 (and results work for Type-F as well).
@@ 3. Minimal product
There have been many specific situations where minimal product has been employed. For all $L_{i}$ being full spheres, the structure produces Clifford type minimal surfaces. When these $L_{i}$ are R-spaces, [OS] studied whether cones over minimal products of them are area-minimizing. For general minimal submanifolds as inputs, see [X] for a proof through the Takahashi Theorem. Around 2016 it was rediscovered independently by [TZ] and [CH] using different methods. Reader may choose either familiar way to understand the structure.
3.1. Formula of $\alpha^{2}$ for minimal product.
For simplicity, we mainly focus on $n=2$ in the paper and all related results for $n\geq 3$ can be gained by induction. In [TZ], information about the second fundamental form of minimal product has been clearly figured out. Let $k_{1},k_{2}$ be the dimension of $L_{1},L_{2}$ respectively. Take $x_{1}\in L_{1}$ and $x_{2}\in L_{2}$. Together with $\eta_{0}=(\lambda_{2}x_{1},-\lambda_{1}x_{2})$, orthonormal bases $\{\sigma_{1},\cdots,\sigma_{N_{1}-k_{1}}\}$ and $\{\tau_{1},\cdots,\tau_{N_{2}-k_{2}}\}$ of $T_{x_{1}}^{\perp}L_{1}$ and $T_{x_{2}}^{\perp}L_{2}$ lead to an orthonormal basis
@@
3.2. Formula of normal radius for minimal product
Another vital quantity in Lawlor’s criterion is the normal radius, so in this subsection we derive its formula rigorously for the minimal product structure.
Let $R_{1},\,R_{2},\,R$ be the normal radii of $L_{1},\ L_{2}$ and $L_{1}\times L_{2}$ in corresponding Euclidean spheres respectively. For $(\lambda_{1}x_{1},\lambda_{2}x_{2})$ in $L_{1}\times L_{2}$, the combination
@@ 3.2. Formula of normal radius for minimal product
where $\xi_{1}$, $\xi_{2}$ are unit normals of $L_{1}$ and $L_{2}$ at $x_{1}$ and $x_{2}$ respectively, and $a_{1}^{2}+a_{2}^{2}+a_{0}^{2}=1$, can exhaust all unit normals of $L_{1}\times L_{2}$ at $(\lambda_{1}x_{1},\lambda_{2}x_{2})$. Then, in $\mathbb{S}^{N_{1}+N_{2}+1}$, each unit normal induces a normal geodesic
@@
The formula of normal radius for minimal product can be derived.
Lemma 3.2.
Remark 3.3.
Note that the result is in fact valid when $\lambda_{1}$ and $\lambda_{2}$ are just two nonzero real numbers $a_{1},a_{2}$ with $a_{1}^{2}+a_{2}^{2}=1$. However, we will restrict ourselves in the minimal product construction unless otherwise specified.
Remark 3.4.
If some $L_{i}$ is the full sphere $\mathbb{S}^{N_{i}}$, then $R_{i}$ is defined to be $\pi$.
Proof.
The three possibilities for $R$ based on (3.4) and (3.5) are the followings.
@@
So, to search for the value of $R$ one should focus on Case II to see whether
(3.9) $\cos R=\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}.$
can be reached. Note that it is always true that
@@
Hence $\sin R\neq 0$. So, in Case II, $a_{2}$ has to be zero in (3.5). Now we want to solve:
Given $\phi_{1}=R_{1}$ for some $x_{1}$ and direction $\xi_{1}$ realizing normal radius $R_{1}$ of $L_{1}$ in $\mathbb{S}^{N_{1}}$, can we get $a_{0},\,a_{1}$ (with $a_{0}^{2}+a_{1}^{2}=1$) and $\nu=a_{1}(\xi_{1},0)+a_{0}\eta_{0}$ to make (3.4), (3.5) and (3.9) hold?
This can be done by
@@
Therefore $\cos R$ takes the larger value of $\lambda_{1}^{2}\cos R_{1}+\lambda_{2}^{2}$ and $\lambda_{1}^{2}+\lambda_{2}^{2}\cos R_{2}$ and (3.6) follows for genuine submanifold inputs.
Regarding Remark 3.4, when $L_{1}$ (or $L_{2}$ or both) is a full sphere, the normal radius $R_{1}$ (or $R_{2}$ or both) cannot be well defined as in Definition 2.1. For our purpose, we always define the normal radius of a full sphere $\mathbb{S}^{N_{1}}$ in itself to be $\pi$. The reason is the following. It can be realized as a totally geodesic sphere in $\mathbb{S}^{N_{1}+\ell}$ for $\ell>0$ and now $\mathbb{S}^{N_{1}}$ has normal radius $\pi$. So the above argument works for the minimal product and the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ can be completely dropped as they contribute nothing (let $\xi$ be a unit normal vector to the minimal product at some point and if $\xi$ has a nonzero part coming from the last $\ell$ components of $\mathbb{R}^{N_{1}+1+\ell}$ then the geodesic with initial velocity vector $\xi$ in $\mathbb{S}^{N_{1}+\ell+N_{2}}$ will run at least length $\pi$ to return to the minimal product and hence this will not realize the normal radius as the normal radius is strictly less than $\pi$ for the minimal product - not a totally geodesic sphere, actually no more than $\frac{\pi}{2}$ by Lemma 4.6).  [^3] For example, the normal radius of minimal product $\mathbb{S}^{N_{1}}\times\mathbb{S}^{N_{2}}$ in $\mathbb{S}^{N_{1}+N_{2}+1}$ is realized by $\eta_{0}$ and can reach $\frac{\pi}{2}$ only when $N_{1}=N_{2}$ according to (3.9), cf. the beautiful uniform formula in [Law] for normal radius of minimal product among full spheres. Thus, (3.6) still holds with the understanding in Remark 3.4. ∎
@@
Similar statement formulated independently and some partial discussion included in [JCX].